OpenStudy (anonymous):

Find the circumference and area of a circle with a 12-cm diameter.

OpenStudy (anonymous):

where do i even start

OpenStudy (anonymous):

Do you know these formulas? C=pi*d A=pi*r^2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

please explain

OpenStudy (anonymous):

haha, that explains it then. C is circumference and A is area. d is diameter, which is 2r (or 2 times the radius) and r is radius. pi is\[\pi \]

OpenStudy (anonymous):

So, you can plug 12 into d for C=pi*d. And you can divide 12 by 2 to find r, which you can plug into A=pi*r^2 Does it make sense?

OpenStudy (anonymous):

area or circumference. which is the radius?

OpenStudy (anonymous):

radius is r. hmm, try the first equation for circumference: C=pi*d and I can coach from there, if you want.

OpenStudy (anonymous):

c=pi*12 that is how it would be set up right?

OpenStudy (anonymous):

yeah. do you know if you're supposed to estimate pi as 3.14?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ok. then just leave the answer as 12*pi. How far can you get with the second equation?

OpenStudy (anonymous):

a=pi*12

OpenStudy (anonymous):

a=3.14*12= so area is 37.68

OpenStudy (anonymous):

well, actually, \[a=\pi r^{2}\] so you would divide 12 by 2 to get r=6. Then square 6 to get 36, and multiply that by pi

OpenStudy (anonymous):

so the frmula would look like a=3.14*6Squared

OpenStudy (anonymous):

yeah. I'm not sure if you knew but "^" stands for "to the power of", btw.

OpenStudy (anonymous):

oh ok now i knw

OpenStudy (anonymous):

thank you so much for your help

OpenStudy (anonymous):

no problem :)